I need help solving this problem
2x + 2y - 4z + w = -5
3y + 4z - w = 0
2z -7w = -6
5w = 15

Answer by richwmiller(17219)

5w=15
w=3
2z-7w=-6
2z-7*3=-6
2z-21=-6
2z=15
z=7.5
3y+4z-w=0
3y+4*7.5-3=0
3y+30-3=0
3y=-27
y=-9
2x+2y-4z+w=-5
2x-18-30+3=-5
2x-45=-5
2x=40
x=20